40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Feb 18 12:23:09 2012, Den, Update, IOO, MC  mc_trend.png
    Reply  Sat Feb 18 16:58:59 2012, kiwamu, Update, IOO, MC suspension realigned 
Message ID: 6295     Entry time: Sat Feb 18 16:58:59 2012     In reply to: 6294
Author: kiwamu 
Type: Update 
Category: IOO 
Subject: MC suspension realigned 

[ Den / Kiwamu]

 We have realigned the MC suspensions so that the WFS servos are smoothly engaged.

Now it seems working fine. The beam pointing to the interferometer also looks okay.

The WFSs control kept failing to engage the servos because of large misalignments in the MC suspensions.

When the TEM00 was locked, the transmitted light was only about 1200 counts and the reflected light was about 2.8 counts.

We tweaked MC1, MC2 and MC3.

Quote from #6294

When I came to the 40m this afternoon, the MC was unlocked. Here is the trend of MC_F for last 2 hours

ELOG V3.1.3-